differentiate the following function:
f(x)=(x^2+1/x^3 )(x^3-x^2+1)

Answer by Cromlix(4381)

Hi there,
Straighten out first.
f(x)=(x^2+1/x^3 )(x^3-x^2+1)
f(x) = (x^2 + x^-3)(x^3-x^2+1)
Multiply next
f(x) = x^5 - x^4 + x^2 + 1 - x^-1 + x^-3
f'(x) = 5x^4 - 4x^3 + 2x + x^-2 - 3x^-4
f'(x) = 5x^4 - 4x^3 + 2x + 1/x^2 - 3/x^4
Hope this helps :-)